TUN A battery is rated that it provides 4.0mW of power at 7.0V when tully chargeu. 0.76 mA 1.8 kA 12 kA O 0.57 mA
Answer #1

here,

the power of battery , P = 4 mW = 0.004 W

the voltage of battery , V = 7 V

the current in the circuit , I = P /V

I = 0.004 /7 A = 0.57 * 10^-3 A

I = 0.57 mA